OpenStudy (jhannybean):

Calculus 3: simplifying expression

OpenStudy (anonymous):

\[ \sin^2x+\cos^2x=1 \]

OpenStudy (jhannybean):

\[\int_{0}^{\pi}\int_{0}^{a\cos(\theta)} \frac{a}{\sqrt{a^2-r^2}}r dr d\theta\]

OpenStudy (jhannybean):

pulled out the a, ofc.

OpenStudy (dan815):

oh this one is different trig identity

OpenStudy (anonymous):

Wait a moment, did it get subbed in the limit already?

OpenStudy (dan815):

oh nvm i didnt see the r there

OpenStudy (dan815):